# Differentially Private Matchings

### v2

@@ 8 Arbitrary Edge-Order Continual Release Implicit Matchings
Input: Arbitrary edge-order stream $S$, privacy parameter $\varepsilon>0$, approximation parameter $\eta\in(0,1]$,
@@ 3 $c\leftarrow a\eta^{-1}\log(n)$
5 for edge $e_{t}\in S$ do
       // $\nu(G_{t})$ denotes the size of the maximum matching of the dynamic graph $G_{t}$ at time $t$.
6       while EstimateSVT.ProcessQuery$(\nu(G_{t}),(1+\eta)^{j_{t}})$ is “above” do
@@ 7             $j_{t}\leftarrow j_{t}+1$
11             $\operatorname{estimate}\leftarrow$ ($\nicefrac{{\varepsilon}}{{3c}}$)-edge DP estimate of the size of the current maximum matching
12      $j_{t+1}\leftarrow j_{t}$
@@ 13       Output solution, estimate
Algorithm 6 Arbitrary Edge-Order Continual Release Edge-DP Matching
@@ 8.1.1 Privacy Proof
Here the notation $<t$ is a shorthand that denotes the event that $(j_{\tau},\operatorname{solution}_{\tau},\operatorname{estimate}_{\tau})_{\tau=1}^{t-1}=(J_{\tau},B_{\tau},M_{\tau})_{\tau=1}^{t-1}$ for $t\geq 2$ and the trivial event of probability $1$ if $t=1$. Thus our goal is to bound the product of ratios
@@
Now, conditioned on the event that $j_{t-1}=J_{t-1}$, it holds that $j_{t}$ is independent of $\operatorname{solution}_{t-1}$ and $\operatorname{estimate}_{t-1}$. Hence the first product is at most $e^{\varepsilon/3}$ by the privacy guarantees of the SVT (Theorem A.3). For the second and third products, remark that given $j_{t}=J_{t-1}=\tilde{j}_{t}$, then $\operatorname{solution}_{t}=B_{t-1}=\widetilde{\operatorname{solution}}_{t}$ with probability $1$ and similarly for $\operatorname{estimate}_{t}$. Since $(J_{t})_{t}$ is a non-decreasing non-negative integer sequence bounded above by $c=a\log(n)/\eta$, at most $c=a\log(n)/\eta$ of the ratios from the second and third products are not 1. We can bound each of these ratios by $e^{\varepsilon/3c}$ using the individual privacy guarantees of Algorithm 2 (Theorem 1.3) and the Laplace mechanism. This concludes the proof. ∎
@@ 8.1.2 Utility Proof
Algorithm 6 outputs a sequence of implicit $b$-matchings with the following guarantee with probability at least $1-1/\operatorname{poly}(n)$: (i) $b=O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)$and (ii) each implicit solution contains a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
@@
(i) follow directly from Theorem 1.3 and the fact that we require each call to Algorithm 2 to satisfy ($\nicefrac{{\varepsilon}}{{3c}}$)-privacy for $c=a\log(n)/\eta$.
To see (ii), we first observe that this certainly holds at timestamps when $\operatorname{solution}$ is updated. By the guarantees of SVT, $(1+\eta)^{j_{t}}$ is a $(1+\eta,O(c\log(n)/\varepsilon))$-approximate estimate of size of the current maximum matching. Hence we incur at most this much error in between updates. ∎
Combining Lemma 8.1 and Lemma 8.2 yields the proof of Theorem 1.11. The pseudocode for Theorem 1.11 is given in Algorithm 6.
8.2 Proof of Theorem 1.12
8.2.1 Extension to Node-Differential Privacy with Public Arboricity Bound
We now describe the necessary modifications to adapt our edge DP continual release algorithm to the challenging node DP setting. Similary to Section 6.2.1, we first present an algorithm assuming a public bound $\tilde{\alpha}$ on the arboricity.
@@ 8.2.1 Extension to Node-Differential Privacy with Public Arboricity Bound
Fix $\eta\in(0,1]$. Given a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the input graph, Algorithm 7 is an $\varepsilon$-node DP algorithm in the arbitrary order continual release model that outputs implicit $b$-matchings. Moreover, with probability at least $1-1/\operatorname{poly}(n)$, (i) $b=O\left(\frac{\tilde{\alpha}\log^{2}(n)}{\eta^{2}\varepsilon}\right)$and (ii) each implicit solution contains a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
@@
Let the sparsified graph $H$ be obtained from the input graph $G$ as described above. Write $c=O(\log(n)/\eta)$ be the maximum SVT budget for “above” queries (Algorithm 10). Our algorithm proceeds as following for each edge: Similar to Algorithm 6 of Algorithm 6, we use an SVT instance to check when to update our solution. Since the maximum matching size sensitivity is 1 for node-neighboring graphs, this step requires no change from the edge-DP algorithm. At timestamps where we must update the solution, we run a static $O(\nicefrac{{\varepsilon}}{{c\Lambda}})$-edge DP implicit matching algorithm (Algorithm 2) on the sparsified graph $H$. This is similar to Algorithm 6 of Algorithm 7 except we need to execute the underlying algorithm with a smaller privacy parameter. Finally, we also estimate the size of the current maximum matching. Once again, this step is unchanged from Algorithm 6 of Algorithm 6. In particular, we note that we only use the sparsified graph $H$ for updating the solution and not for estimating the current maximum matching size.
@@
1 $\Lambda\leftarrow 5(1+\nicefrac{{5}}{{\eta}})2\tilde{\alpha}$
@@ 8             $E[H]\leftarrow E[H]\cup\{e_{t}\}$
      // Check if $\nu(G_{t})$ has significantly increased using SVT with total privacy budget $\nicefrac{{\varepsilon}}{{3}}$ (Algorithm 6 of Algorithm 6)
@@ 10       Output $\operatorname{solution},\operatorname{estimate}$
Algorithm 7 Arbitrary Edge-Order Continual Release Node-DP Matching Algorithm
@@
The privacy proof is identical that in the proof of Theorem 1.11. The utility guarantees follow similarly, with the exception that we use the approximation guarantees of Theorem 6.3 rather than Theorem 1.3. ∎
8.2.2 Removing the Assumption of Public Bound
@@
9 Adjacency-List Order Continual Release Implicit Matchings
In the adjacency-list order insertion streams, nodes arrive in an arbitrary order, and once a node arrives, all edges adjacent to the node arrive in an arbitrary order. Remark that in the adjacency-list order stream, all edges arrive twice, once per endpoint in the stream.
In this section, we give $\varepsilon$-DP algorithms for implicit matching in the continual release model with adjacency-list order streams. Our results for another stream type (arbitrary edge-order streams) are presented in Section 8. The adjacency-list order stream ensures each node that arrives in this model will be followed by its edges where the edges arrive in an arbitrary order. Our algorithm is a straightforward implementation of Algorithm 2 in the continual release model.
@@ 9 Adjacency-List Order Continual Release Implicit Matchings
Below, we restate our final result (Theorem 1.13) about the privacy and utility guarantees of our continual release algorithm. See 1.13
Detailed Algorithm
@@ 37             $K\leftarrow K\cup\{e_{i}\}$
10 Private Implicit $b$-Matchings
In this section, we extend our all of our results to the $(2+\eta)$-approximate $b$-matchings, including our node-DP results, which use a novel $b$-matching arboricity sparsifier. The key difference between the results in this section and our previous result is that our approximation bounds are given in terms of the maximum $b$-matching given large enough $b$. In particular, we satisfy the $b$-matching constraints (that our matching has degree less than $b$ for all vertices), with high probability, and output a $(2+O(\eta))$-maximum $b$-matching). Thus, our approximation is no longer a bicriteria approximation in this setting.
To obtain our $b$-matching results, we prove a stronger version of our matching result. Namely, we show that we can guarantee that our implicit output is at least the size of a $(2+\eta)$-approximate maximum $b^{\prime}$-matching for some $b^{\prime}<b$. In particular, $b$ has to be an additive factor of $\Omega\left(\frac{\ln(n)}{{\eta^{2}}\varepsilon}\right)$ greater than $b^{\prime}$ to ensure the returned solution is comparable to a $2$-approximate $b^{\prime}$-matching. Furthermore, if $b=\Omega\left(\frac{\ln(n)}{\eta^{3}\varepsilon}\right)$, then we get a $(2+\eta)$-approximate maximum $b$-matching.
10.1 $\varepsilon$-LEDP $b$-Matchings
Our revised algorithm is the same algorithm as Algorithm 2 except with a different lower bound on $b$ depending on what $b^{\prime}$-maximal matching we want to guarantee in our returned solution. This is our only change to the algorithm. When our lower bound on $b$ contains a factor of $1/\eta^{2}$, then we obtain a $(2+\eta)$-approximate maximum $b$-matching. We restate our modified algorithm below with the modified constants highlighted in blue. The constants do not affect our privacy proof; hence, we only give an updated version of our utility analysis following our pseudocode in Algorithm 2. Since the proof and algorithm only require minor changes for the generalization, we replicate our previous algorithm and proof and highlight our changes in blue.
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$, $b^{\prime}>0$ parameter for $b^{\prime}$-matching, matching parameter $b\geq{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}}$, constant $\eta\in(0,1)$, and constant $c\geq 3$
4
5for each node $u\in V$ do
7       $M(u)\leftarrow\varnothing$
8       $A(u)\leftarrow\infty$
9
10for iteration $i=1$ to $n$, let $v=v_{i}$ do
12             $\nu_{i}(u)\leftarrow\text{Lap}(8/\varepsilon^{\prime})$
13             if $|M(u)|+\nu_{i}(u)\geq\tilde{b}(u)$  then
14                   $u$ releases $A(u)\leftarrow i$
15
16
17       if $A(v)>i$ ($v$ has not satisfied its matching condition)  then
18             for subgraph index $r=0,\ldots,\lceil\log_{1+\eta}(n)\rceil$  do
20                   $|\widetilde{W}_{r}(v)|=|W_{r}(v)|+\text{Lap}(2/\varepsilon^{\prime})$
21
22            $|\widetilde{M}_{i}(v)|=|M(v)|+\text{Lap}(2/\varepsilon^{\prime})$
24             $v$ releases $r_{v}$
25             for $u\in W_{r_{v}}(v)$ do
26                   $M(u)\leftarrow M(u)\cup\{v\}$
27                   $M(v)\leftarrow M(v)\cup\{u\}$
28
29
Algorithm 9 $\varepsilon$-LEDP Approximate Maximum $b$-Matching
We now restate and prove Theorem 1.4. See 1.4
The proof is implied by the following lemma which is a restatement with detailed constants.
Lemma 10.1.
Proof.
First, all of the Laplace noise (Algorithms 9, 9, 9 and 9) are drawn from distributions with expectation 00. Let $X$ be a random variable drawn from one of these Laplace distributions. By Lemma A.1, we have that each of the following holds with probability $1-1/n^{3c}$:
Case I: Suppose that Algorithm 9 does not execute, i.e. $v$ has already satisfied its matching condition as evaluated on Algorithms 9, 9 and 9. Then we know that $|M(v)|+\nu_{i}(v)\geq\tilde{b}(v)$ by definition. We then have that
$|M(v)|$ $\geq b-{60}c\ln(n)/\varepsilon^{\prime}$
$\geq{b}-\frac{72c\ln(n)}{\varepsilon^{\prime}}$
$\geq{b}-\frac{576c\ln(n)}{\eta\varepsilon}$
$>{b^{\prime}}\,.$ (by $b\geq b^{\prime}+\frac{576c\ln(n)}{\eta\varepsilon}$)
$\leq\frac{238c\ln(n)}{\eta^{2}\varepsilon}+{b^{\prime}}$
$\leq b$
and all of $W_{0}(v)$ is matched.
$|M(v)|+|W_{r}(v)|+\frac{24c\ln(n)}{\varepsilon^{\prime}}$ $>b\,.$
In other words,
$|M(v)|$ $>b-|W_{r}(v)|-\frac{24c\ln(n)}{\varepsilon^{\prime}}$
$|M(v)|$ $>b-\frac{162c\ln(n)}{\eta^{2}\varepsilon}$
In other words, $v$ was already matched from a previous iteration.
with probability $1-\frac{1}{n^{4c}}$.
$\geq b^{\prime}-|M(v)|\,.$
$|M(u)|$ $<\tilde{b}(u)-\nu_{i}(u)$
$|M(u)|$ $<b.$
$=|M(v)|+|W_{r}(v)|.$
Thus in this case, $v$ is also matched to at most $b$ neighbors, as desired. ∎
We now restate and prove Theorem 1.5. See 1.5 The proof follows from the following corollary of Lemma 10.1.
Corollary 10.2 ($(2+\eta)$-Approximate Maximum $b$-Matching).
Before proceeding with the proof, we recall that $b$-matchings have the following integral relaxation [schrijver2003combinatorial, Theorem 33.2], where $\partial(U)$ denotes the set of outgoing edges from $U\subseteq V$.
$\sum_{uv\in E}x(uv)$ $\leq b$ $\forall v\in V$
$x$ $\in[0,1]^{E}$
Proof.
Choose $b^{\prime}$ such that
10.2 Node-Differentially Private $b$-Matchings
Similar to Section 6, we can implement our implicit $b$-matching algorithm in node-DP setting with the help of an arboricity sparsifier. However, this requires extending Solomon’s sparsifier [solomon2021local] to the setting of $b$-matchings. Doing so and repeating the steps in Section 6 yields a proof of Theorem 1.9. See 1.9
The rest of this section serves to prove the necessary sparsifier extension.
10.2.1 Bounded Arboricity Sparsifiers for $b$-Matchings
Theorem 10.3.
$\mu_{b}(H)\leq\mu_{b}(G)\leq(1+\eta)\mu_{b}(H).$
[…]
Proof.
We argue by induction on $b$. The base case of $b=1$ is guaranteed by Theorem 6.2. Suppose inductively the statement holds for $b-1$ for some $b\geq 2$ so that our goal is to extend it to the case $b$.
Remark that $M_{1}^{\prime}\cup M_{2}\cup\dots\cup M_{b}$ is a $b$-matching within $G$ by construction. Consider $G_{-1}\coloneqq G-M_{1}^{\prime}$, so that $M_{2}\cup\dots\cup M_{b}$ is a $(b-1)$-matching within $G_{-1}$. We claim that $E_{H}\cap E(G_{-1})$ contains a $(b-1)$-matching of size at least $\frac{1}{1+\eta}|M_{2}\cup\dots\cup M_{b}|$. To prove this claim, we will rely on the induction hypothesis. Define
Then each $T_{-1}(v)$ contains at least $\min(d_{G_{-1}}(v),\Lambda_{b-1})$ neighbors of $v$ in $G_{-1}$. Furthermore, $E_{-1}\subseteq(E_{H}\setminus M_{1}^{\prime})$ by construction. Thus by the induction hypothesis, $E_{-1}$ and hence $H$ contains a $(1+\eta)$-approximate maximum $(b-1)$-matching $N$ of $G_{-1}$.
We have identified disjoint edge subsets $M_{1}^{\prime},N\subseteq E_{H}$ such that $M_{1}^{\prime}$ is a matching in $G$ with size at least $\frac{1}{1+\eta}|M_{1}|$ and $N$ is a $(b-1)$-matching in $G-M_{1}^{\prime}$ with size at least $\frac{1}{1+\eta}|M_{2}\cup\dots\cup M_{b}|$. Thus $M_{1}^{\prime}\cup N$ is a $b$-matching in $G$ with size at least
10.3 Arbitrary Edge-Order Continual Release $b$-Matchings
Similarly, we can extend our $b$-matching algorithms to the continual release setting. Here we state the guarantees for an $\varepsilon$-edge DP algorithm for arbitrary edge-order stream, but our algorithms also extend to the node-DP setting.
Theorem 10.4.
For $\eta\in(0,1)$ and $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\Omega\left(\frac{\log^{2}(n)}{\eta^{3}\varepsilon}\right),$ there is an $\varepsilon$-edge DP algorithm that outputs a sequence of implicit $b$-matchings in the billboard model such that each $b$-matching is a $(2+\eta)$-approximation with respect to the maximum $b^{\prime}$-matching size, with probability at least $1-\frac{1}{n^{c}}$.
11 Other Lower Bounds
11.1 Lower Bound for Implicit Matchings
We now give our lower bound for implicit solutions, which essentially matches our upper bound from Theorem 1.3. While it does not apply to all algorithms in the billboard model, it applies to the implicit solutions that our algorithms use (Definition 2.6).
Theorem 11.1.
Note that this essentially matches Theorem 1.3 (up to constants) by setting $\beta=1/\mathrm{poly}(n)$. In the rest of this section we will prove Theorem 11.1. Recall that, given an input graph $H$, our implicit solutions essentially output a subset of vertices for each node $x$ whose intersection with $x$’s private adjacency list gives the nodes $x$ is matched to. We denote these public subsets of vertices by $\mathcal{S}=\{S_{x}\}_{x\in V}$ where $S_{x}$ is the public subset of vertices for node $x$. That is, $\mathcal{S}=\{S_{x}\}_{x\in V}$ is the implicit solution generated by $\mathcal{A}$. We say that $\mathcal{A}$ includes an edge $\{u,v\}$ if either $u\in S_{v}$ or $v\in S_{u}$. In other words, $\mathcal{A}$ includes an edge if that edge is an edge in $H(\mathcal{S})$.
Proof of Theorem 11.1.
Pr[number of non-edges from r to T in $\Pr\left[\text{number of non-edges from $r$ to $T$ in }H(\mathcal{S})\geq\frac{|T|}{2}\right]\leq\frac{2e^{2\varepsilon}\beta|T|}{|T|}=2e^{2\varepsilon}\beta.$
11.2 Lower Bound for Node DP Matching Sparsifiers
Theorem 11.2.
Let $\mathcal{A}$ be an $(\varepsilon,\delta)$-node DP algorithm which, given an input graph $G=(V,E)$, outputs an implicit matching sparsifier $E^{\prime}$. Then if $\nu(G)\leq(1+\eta)\cdot\text{E}[\nu(E^{\prime}\cap E)]$, the maximum degree of $E^{\prime}\cap E$ is at least $\left(\frac{1}{e^{\varepsilon}(1+\eta)}-\delta\right)(n-1)$ even if $G$ has arboricity $1$.
Proof.
Fix $V$ with $|V|=n$, and let $r\in V$. For any $S\subseteq V\setminus\{r\}$, let $G_{S}$ be the graph whose edge set consists of an edge from $r$ to all nodes in $S$ (i.e., a star from $r$ to $S$ and all other nodes having degree 00). Note that all such graphs have arboricity $1$. A similar symmetry argument to the other lower bounds implies that without loss of generality, we may assume that the probability that $\mathcal{A}$ includes some edge $\{r,v\}$ is the same for all $v\in S$ (we will denote this probability by $p_{S}$).
Fix some node $v\in S\setminus\{r\}$, and consider what happens when we run $\mathcal{A}$ on the graph $G_{\{v\}}$. Since $\text{E}[\nu(E^{\prime}\cap E)]\geq\frac{\nu(G_{\{v\}})}{1+\eta}$ by assumption, we know that $p_{\{v\}}=\Pr[\{r,v\}\in E^{\prime}]\geq\frac{1}{1+\eta}$.
[…]

### v3

@@ 9.1 Edge-DP Arbitrary Edge-Order Continual Release Implicit $b$-Matchings
Input: Arbitrary edge-order stream $S$, privacy parameter $\varepsilon>0$, approximation parameter $\varrho\in(0,1]$,
@@ 3 $c\leftarrow\log_{1+\varrho}(n)$
5
61exfor edge $e_{t}\in S$ do
7
  1ex// $\nu(G_{t})$ denotes the size of the maximum matching of the dynamic graph $G_{t}$ at time $t$.
[…]
@@ 9    $j_{t}\leftarrow j_{t}+1$
14    $\operatorname{estimate}\leftarrow$ ($\nicefrac{{\varepsilon}}{{3c}}$)-edge DP estimate of the size of the current maximum $b^{\prime}$-matching
15
[…]
@@ 17   Output solution, estimate
Algorithm 6 Arbitrary Edge-Order Continual Release Edge-DP Implicit $b$-Matching
@@ 9.1.1 Privacy Proof
Here the notation <t<t is a shorthand that denotes the event that $(j_{\tau},\operatorname{solution}_{\tau},\operatorname{estimate}_{\tau})_{\tau=1}^{t-1}=(J_{\tau},B_{\tau},M_{\tau})_{\tau=1}^{t-1}$ for $t\geq 2$ and the trivial event of probability $1$ if $t=1$. Thus our goal is to bound the product of ratios
@@
Now, conditioned on the event that $j_{t-1}=J_{t-1}$, it holds that $j_{t}$ is independent of $\operatorname{solution}_{t-1}$ and $\operatorname{estimate}_{t-1}$. Hence the first product is at most $e^{\varepsilon/3}$ by the privacy guarantees of the SVT (Theorem A.3). For the second and third products, remark that given $j_{t}=J_{t-1}=\tilde{j}_{t}$, then $\operatorname{solution}_{t}=B_{t-1}=\widetilde{\operatorname{solution}}_{t}$ with probability $1$ and similarly for $\operatorname{estimate}_{t}$. Since $(J_{t})_{t}$ is a non-decreasing non-negative integer sequence bounded above by $c=a\log(n)/\eta$, at most $c=a\log(n)/\eta$ of the ratios from the second and third products are not 1. We can bound each of these ratios by $e^{\varepsilon/3c}$ using the individual privacy guarantees of Algorithm 2 (Corollary 5.1) and the Laplace mechanism. This concludes the proof. ∎
@@ 9.1.2 Utility Proof
Let $\varepsilon\in(0,1)$, $\varrho,\eta\in(0,1)$, and $b\geq\frac{(1+\eta)^{2}}{1-\eta}\cdot b^{\prime}+O\left(\frac{\log^{2}(n)}{\varrho\eta^{2}\varepsilon}\right)$. Algorithm 6 outputs a sequence of implicit $b$-matchings with the following guarantee with probability at least $1-1/\operatorname{poly}(n)$: each implicit solution contains a $\left(2+\varrho,O\left(\frac{\log^{2}(n)}{\varrho\varepsilon}\right)\right)$-approximate maximum matching.
@@
The bound on $b$ follow directly from Theorem 2.2 and the fact that we require each call to Algorithm 2 to satisfy ($\nicefrac{{\varepsilon}}{{3c}}$)-privacy for $c=\log_{1+\varrho}(n)$.
To see the approximation guarantee, we first observe that this certainly holds at timestamps when $\operatorname{solution}$ is updated. By the guarantees of SVT, $(1+\varrho)^{j_{t}}$ is a $(1+\varrho,O(\log^{2}(n)/\varepsilon\varrho))$-approximate estimate of size of the current maximum matching. Hence we incur at most this much error in between updates. ∎
Combining Lemma 9.4 and Lemma 9.5 yields the proof of Corollary 9.1. The pseudocode for Corollary 9.1 is given in Algorithm 6.
9.1.3 Corollaries
As a special case of our general edge-DP algorithm (Theorem 2.9) when $\eta=\nicefrac{{1}}{{2}}$ and $b^{\prime}=1$, we recover the guarantees stated in Corollary 9.1.
In order to obtain a non-bicriteria approximation when $b=\Omega\left(\frac{\log^{2}(n)}{\eta^{4}\varepsilon}\right)$, we use the same technique as the static edge-DP algorithm of executing the general algorithm (Theorem 2.9) with $b^{\prime}=\frac{b}{1+O(\eta)}$. The proof is identical to that of Corollary 5.7. This yields a proof of Theorem 9.2.
9.2 Node-DP Arbitrary Edge-Order Continual Release Implicit $b$-Matchings
We now describe the necessary modifications to adapt our edge DP continual release algorithm to the challenging node DP setting. Similary to Section 7.1, we first present an algorithm assuming a public bound $\tilde{\alpha}$ on the arboricity.
@@ 9.2 Node-DP Arbitrary Edge-Order Continual Release Implicit $b$-Matchings
Fix $\eta\in(0,1]$. Given a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the input graph, Algorithm 7 is an $\varepsilon$-node DP algorithm in the arbitrary order continual release model that outputs implicit $b$-matchings. Moreover, with probability at least $1-1/\operatorname{poly}(n)$, (i) $b=O\left(\frac{\tilde{\alpha}\log^{2}(n)}{\eta^{2}\varepsilon}+\frac{b^{\prime}\log^{2}(n)}{\eta\varepsilon}\right)$and (ii) each implicit solution contains a $\left(2+\eta,O\left(\frac{b^{\prime}\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
@@
Let the sparsified graph $H$ be obtained from the input graph $G$ as described above. Write $c=O(\log(n)/\eta)$ be the maximum SVT budget for “above” queries (Algorithm 9). Our algorithm proceeds as following for each edge: Similar to Algorithm 6 of Algorithm 6, we use an SVT instance to check when to update our solution. The only change is that $b^{\prime}$-matching sensitivity is $b^{\prime}$ for node-neighboring graphs. At timestamps where we must update the solution, we run a static $O(\nicefrac{{\varepsilon}}{{c\Lambda}})$-edge DP implicit matching algorithm (Algorithm 2) on the sparsified graph $H$. This is similar to Algorithm 6 of Algorithm 7 except we need to execute the underlying algorithm with a smaller privacy parameter. Finally, we also estimate the size of the current maximum matching. Once again, this step is unchanged from Algorithm 6 of Algorithm 6, with the slight exception that the sensitivity is now $b^{\prime}$. In particular, we note that we only use the sparsified graph $H$ for updating the solution and not for estimating the current maximum matching size.
@@
1 $\Lambda\leftarrow 5(1+\nicefrac{{5}}{{\eta}})2\tilde{\alpha}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}+(b^{\prime}-1)}$
@@ 8    $E[H]\leftarrow E[H]\cup\{e_{t}\}$
10
  1ex// Check if $\nu(G_{t})$ has significantly increased using SVT with sensitivity $b^{\prime}$ and total privacy budget $\nicefrac{{\varepsilon}}{{3}}$ (Algorithm 6 of Algorithm 6)
@@ 11   Output $\operatorname{solution},\operatorname{estimate}$
Algorithm 7 Arbitrary Edge-Order Continual Release Node-DP $b$-Matching Algorithm
@@
The privacy proof is identical to that of Theorem 2.9. The utility guarantees follow similarly, with the exception that we use the approximation guarantees of Theorem 2.6 rather than Theorem 2.2. ∎
Public Arboricity Bound.
@@
9.3 Adjacency-List Order Continual Release Implicit Matchings
In this section, we give $\varepsilon$-DP algorithms for implicit matching in the continual release model with adjacency-list order streams. Our results for another stream type (arbitrary edge-order streams) are presented in Section 9.1. The adjacency-list order stream ensures each node that arrives in this model will be followed by its edges where the edges arrive in an arbitrary order. Our algorithm is a straightforward implementation of Algorithm 2 in the continual release model.
@@ 9.3 Adjacency-List Order Continual Release Implicit Matchings
Detailed Algorithm.
@@ 37    $K\leftarrow K\cup\{e_{i}\}$
Acknowledgments
Felix Zhou acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). Quanquan C. Liu and Felix Zhou are supported by a Google Academic Research Award and NSF Grant #CCF-2453323.